OpenStudy (cheesecakekitten):

If p varies directly with q, and p = 10 when q = 2, what is the value of p when q = 20? A. p = 1 B. p = 40 C. p = 100 D. p = 28

OpenStudy (cheesecakekitten):

@igreen @answermyquestions @lordhelix8th @BloomLocke367

OpenStudy (bloomlocke367):

you can write a proportion.

OpenStudy (bloomlocke367):

\(\Huge\frac{10}{2}=\frac{?}{20}\)

OpenStudy (lordhelix8th):

yea we need the poportion

OpenStudy (solomonzelman):

direct variation, has the equation: \(\large\color{black}{ \displaystyle y=\color{orangered}{\rm k}x }\) In this case, though, your variables are \(p\) and \(q\). \(\large\color{black}{ \displaystyle p=\color{orangered}{\rm k}q }\) you are given that \(p=10\) when \(q=2\) AND you can use this information to solve for \(\large\color{black}{ \displaystyle \color{orangered}{\rm k} }\).

OpenStudy (bloomlocke367):

that works too ^^

OpenStudy (solomonzelman):

your is definitely shorter:)

OpenStudy (cheesecakekitten):

10 / 20 = 0.5

OpenStudy (solomonzelman):

cross multiply to solve for the question mark (which you can denote as x)

OpenStudy (cheesecakekitten):

i used bloomlockes

OpenStudy (solomonzelman):

i figured:)

OpenStudy (bloomlocke367):

you have to multiply \(\Huge 10\times 20\) then divide that by \(\Huge 2\)

OpenStudy (cheesecakekitten):

100

OpenStudy (bloomlocke367):

\(\Huge\color{lime}{Correct~\checkmark}\)

OpenStudy (cheesecakekitten):

This glitch is so annoying i have to leave or reset the page everytime i post something

OpenStudy (bloomlocke367):

Make a post about it in OpenStudy Feedback